The gases in a hair spray can are at a temperature

of 27 °C and a pressure of 207 kPa. If the gases inside

the can reach 400 kPa, the can will explode. At what

temperature would the can explode?


1
Expert's answer
2021-05-11T05:52:49-0400

Answer

Using Gay-Lussac's law


"\frac{P_1}{T_1}=\frac{P_2}{T_2}"


"T_2=T_1\frac{P_2}{P_1}"


"T_1=27+273=300K \\
P_1=207kPa \\
P_2=400 kPa"

Putting the values

We got


"T_2= {300×400\over 207}"

"T_2=579.7 K"

"\implies \\
T_2 = 306.7 \ ^o C" answer